# Latent Logic Tree Extraction for Event Sequence Explanation from LLMs

### v1

@@ 1 Introduction
However, the heterogeneity observed in each patient’s data poses a challenge – each event sequence may exhibit diverse nature of medical histories, treatments, and conditions (Henrich & McElreath, 2003; Laland, 2004). Generating the most relevant and accurate knowledge to explain such heterogeneous data requires sophisticated methods.
@@
Recently, Large Language Models (LLMs) have demonstrated promising human-like reasoning ability as few-shot learners (Brown et al., 2020). When prompted with step-wise explanations of reasoning, these models excel in logical reasoning (Pan et al., 2023), abstract pattern induction (Webb et al., 2023), and social learning (Leng & Yuan, 2023). Despite their success in text-based reasoning tasks, LLMs still face challenges in extending their reasoning capabilities to tabular data (Hegselmann et al., 2023) and discrete event sequences (Shi et al., 2023).
@@
(ii) In the E-step, we use GFlowNets to fine-tune LLMs that enables diverse logic tree generation, which better tackles the heterogeneity issue exhibited for event sequences;
@@ 2.2 Knowledge-Driven Probabilistic Models for Event Sequences
Temporal point process (TPP) provides an elegant probabilistic model for continous-time event sequences, which is characterized by an intensity function. The intensity function represents the occurrence rate of events, which is usually modeled as parametric, nonparameteric, or deep neural network forms. Traditional parametric TPP models like the Hawkes process offer interpretability, but their simplicity limits flexibility. On the other hand, neural-based models, such as RMTPP (Du et al., 2016) and Transformer Hawkes (Zuo et al., 2020), provide expressiveness but are often criticized for their black-box nature and hinder their applications in high-stakes scenarios. In this paper, we aim to generate logic trees from a fine-tuned LLM to inform the functional form of the intensity, which strikes the balance between model flexibility and interpretability. The modeling idea takes inspiration from TL-PP (Li et al., 2020) , as detailed below.
@@
where $f$ is a path from the symbolic logic tree $\mathcal{R}$, $\phi_{f}(\mathbf{X}_{t})$ is the logic-informed feature derived from the number of ordered event combinations in event history $\mathbf{X}_{t}$ satisfying path $f$ (with more details can be found in  (Li et al., 2020)), and $w_{f}$ is the weight corresponding to rule $f$. Given this probabilistic model, we can use the negative log-likelihood of the temporal point process as a loss function to jointly learn weights $w$ and symbolic structure $\mathcal{R}$. Given a sequence $\mathbf{X}$ observed over an interval $[t_{1},t_{L}]$, the negative log-likelihood function is expressed as:
@@
where each $t_{j}$ is the event time and $\mathbf{X}_{j}$ refers to the event sequences up to $t_{j}$. Nevertheless, this learning problem is challenging because it requires learning the parameters $w$ in a continuous space as well as the symbolic structure $\mathcal{R}$ in a discrete space.
@@ 3 Our Proposed LaTee using LLM prior
Instead of eliciting linear thoughts from LLMs, our focus is to extract and reweight symbolic logic trees generated from LLMs to explain the dynamics of the observed event sequences. We hope the obtained logic trees will not only offer personalized explanations for each event sequence but also enable the accurate future events prediction.
@@ 3.1 LLM-Symbolic Integration by Latent Variables
Given event sequences $X$ and next event type $Y$, where $X$ records explicit event sequences with $k$ events $X=\{(t_{i},e_{i})\}_{i=1}^{k}$, $e_{i}$ is the event type for $i$-th event, and $Y=e_{k+1}$ is the next event type after $k$-th event, we are interested in finding a collection of latent symbolic logic trees $\mathcal{R}$, which are composed of various event types that trigger subsequent events and best explain the likelihood of the observed event sequence:
@@
For this mixture latent variable model (LVM), we treat $\mathcal{R}$ as latent variable, $p(\mathcal{R})$ is the prior distribution for the latent logic tress, and the joint likelihood of the event sequences $p(X,Y|\mathcal{R})$ can be derived from temporal point process framework, as shown in Eq. (5).
We will employ LLMs as the prior $p(\mathcal{R})$ for logic trees. Additionally, if we aim to leverage the powerful reasoning and generation capabilities of LLMs to predict $Y$–for instance, in the context of symptom-treatment pairs or question-answer pairs for $(X,Y)$–it becomes intriguing to explore the recommendations of $Y$ given $X$ provided by LLMs. Consequently, we further decompose the mixture LVM (as shown in Eq. (6)) as:
@@
$=\sum_{\mathcal{R}}p_{w}(X|\mathcal{R})p_{LM}(Y|X,\mathcal{R})p_{LM}(\mathcal{R};\phi).$ (8)
We aim to jointly optimize the event likelihood parameter $w$ and the tunable parameters $\phi$ in the prior language model. The challenge in learning arises from the latent variables $\mathcal{R}$. Fortunately, the EM algorithm provides an effective tool for learning mixture models with latent variables. However, in the E-step, we need to analytically evaluate the current posterior $p(\mathcal{R}|X,Y)\propto p_{w}(X|\mathcal{R})p_{LM}(Y|X,\mathcal{R})p_{LM}(\mathcal{R})$, which is intractable due to that the partition function requires the summation over the discrete space of $\mathcal{R}$. To tackle this intractability, variational-EM algorithm (Dempster et al., 1977; Beal, 2003; Koller & Friedman, 2009) can be used to approximate the posterior by optimization. We will address this issue by introducing an amortized EM, where in the E-step we learn GFlowNets to sample from $p(\mathcal{R}|X,Y)$ without the need to calcuate the partition function.
@@ 3.2 Amortized EM framework for Logic Tree Inference
Specifically, in the E-step, we will draw samples from the posterior over the latent symbolic logic tree, denoted as $p_{LM}(\mathcal{R}|X,Y)$, which comes from an amortized sampler of $\mathcal{R}$ with an LLM as its policy. In the M-step, we maximize the log-likelihood of the joint probability of the sampled latent variables $\mathbb{E}_{\mathcal{R}\sim p(\mathcal{R}|X,Y)}[\log p_{w}(X|\mathcal{R})p_{LM}(Y|X,\mathcal{R})p_{LM}(\mathcal{R})]$ with respect to the parameters of $w$. This combination of amortized inference (learning to sample the symbolic logic tree from the language model) and supervised learning (optimizing the likelihood model with the ‘supervision’ involving $\mathcal{R}$ sampled from the amortized posterior) is present in Fig. 2. We illustrate them in detail in below sections.
@@
$\log q_{GFN}(\mathcal{R}_{t+1}|\mathcal{R}_{t}):=\sum_{i=1}^{m}\sum_{k=1}^{W}\log q_{LM}(z_{j+1}^{(i),k}|z_{0}^{(i)}\cdots z_{j}^{(i)}),$ (9)
@@
where $q_{\theta}$ is the conditional GFlowNet policy initialized by a language model $p_{LM}$ conditioned on prefix $X$ and $Y$. In practice, this loss can be minimized by gradient descent on $\theta$ sampled either on-policy or off-policy, just as in reinforcement learning. To predict the event type $Y$ for an unseen event sequence $X$, one can draw samples of $\mathcal{R}$ from $q_{\theta}$ followed by sampling from $p_{LM}(Y|X,\mathcal{R})$.
@@
  until some converge condition
@@ 4.1 Experimental Setup
Competitors. In our study, we categorize competitors into three distinct types. The first category includes prompt-based approaches applied to language models, such as $k$-shot Chain-of-Thought (CoT) (Wei et al., 2022) and Tree-of-Thoughts (ToT) (Yao et al., 2023), which are utilized to generate reasoning chains and make prediction of the last event. The second category involves fine-tuning methods for language models, notably supervised fine-tuning (SFT) and Proximal Policy Optimization (PPO) (Schulman et al., 2017) fine-tuning. The final category consists of advanced neural Temporal Point Process (TPP) models specifically designed for event prediction. Within this group, we focus on AttNHP (Yang et al., 2021), an attention-based TPP whose performance is either on par with or superior to the Neural Hawkes Process (NHP) (Mei & Eisner, 2017) and other attention-based models (Xue et al., 2023a). Additionally, we consider PromptTPP (Xue et al., 2023b), a prompting model based on AttNHP (abrreivated as Pt-AttNHP), tailored for processing streaming events with a retrieval memory mechanism.
@@ Acknowledgement
The authors thank the anonymous reviewers for their careful reading of our manuscript and their many insightful comments and suggestions. Shuang Li’s research was in part supported by the National Science and Technology Major Project under grant No. 2022ZD0116004, the NSFC under grant No. 62206236, Shenzhen Science and Technology Program under grant No. JCYJ20210324120011032, Shenzhen Key Lab of Cross-Modal Cognitive Computing under grant No. ZDSYS20230626091302006, and Guangdong Key Lab of Mathematical Foundations for Artificial Intelligence. Zitao Song acknowledges support from National Research Foundation, Singapore under its AI Singapore Program (AISG Award No: AISG2-GC-2023-009).

### v3

@@ 1 Introduction
However, the heterogeneity observed in each patient’s data poses a challenge – each event sequence may exhibit diverse natures of medical histories, treatments, and conditions (Henrich & McElreath, 2003; Laland, 2004). Generating the most relevant and accurate knowledge to explain such heterogeneous data requires sophisticated methods.
@@
Recently, Large Language Models (LLMs) have demonstrated promising human-like reasoning abilities as few-shot learners (Brown et al., 2020). When prompted with step-wise explanations of reasoning, these models excel in logical reasoning (Pan et al., 2023), abstract pattern induction (Webb et al., 2023), and social learning (Leng & Yuan, 2023). Despite their success in text-based reasoning tasks, LLMs still face challenges in extending their reasoning capabilities to tabular data (Hegselmann et al., 2023) and discrete event sequences (Shi et al., 2023).
@@
(ii) In the E-step, we use GFlowNets to fine-tune LLMs and enable diverse logic tree generation, which better tackles the heterogeneity issue exhibited for event sequences;
@@ 2.2 Knowledge-Driven Probabilistic Models for Event Sequences
Temporal point process (TPP) provides an elegant probabilistic model for continuous-time event sequences, which is characterized by an intensity function. The intensity function represents the occurrence rate of events, which is usually modeled as parametric, nonparametric, or deep neural network forms. Traditional parametric TPP models like the Hawkes process offer interpretability, but their simplicity limits flexibility. On the other hand, neural-based models, such as RMTPP (Du et al., 2016) and Transformer Hawkes (Zuo et al., 2020), provide expressiveness but are often criticized for their black-box nature and hinder their applications in high-stakes scenarios. In this paper, we aim to generate logic trees from a fine-tuned LLM to inform the functional form of the intensity, which strikes a balance between model flexibility and interpretability. The modeling idea takes inspiration from TL-PP (Li et al., 2020) , as detailed below.
@@
where $f$ is a valid path from the symbolic logic tree $\mathcal{R}$, $\phi_{f}(\mathbf{X}_{t})$ is the logic-informed feature derived from the number of ordered event combinations in event history $\mathbf{X}_{t}$ satisfying the path $f$ (with more details can be found in  (Li et al., 2020)), and $w_{f}$ is the weight corresponding to rule $f$. Given this probabilistic model, we can use the negative log-likelihood of the temporal point process as a loss function to jointly learn weights $w$ and symbolic structure $\mathcal{R}$. Given a event sequence $\mathbf{X}=\{(t_{i},e_{i})\}_{i=1}^{L}$ observed over an interval $[0,T]$, the negative log-likelihood of $\mathbf{X}$ is expressed as:
$\mathcal{L}_{w,\mathcal{R}}(\mathbf{X})=$ (5)
@@
where each $t_{j}$ is the event trigger time and $\mathbf{X}_{j}$ refers to the event sequences up to $t_{j}$. Nevertheless, this learning problem is challenging because it requires learning the parameters $w$ in a continuous space as well as the symbolic structure $\mathcal{R}$ in a discrete space.
@@ 3 Our Proposed LaTee using LLM prior
Instead of eliciting linear thoughts from LLMs, our focus is to extract and reweight symbolic logic trees generated from LLMs to explain the dynamics of the observed event sequences. We hope the obtained logic trees will not only offer personalized explanations for each event sequence but also enable accurate future events prediction.
@@ 3.1 LLM-Symbolic Integration by Latent Variables
Given event sequences $X$ and next event type $Y$, where $X$ records explicit event sequences with $k$ events $X=\{(t_{i},e_{i})\}_{i=1}^{k}$, where $t_{i}$ is the $i$-th event time, $e_{i}$ is the $i$-th event type, and $Y=e_{k+1}$ is the next event type after the $k$-th event. We are interested in finding a collection of latent symbolic logic trees $\mathcal{R}$, which are composed of various event types that trigger subsequent events and best explain the likelihood of the observed event sequence:
@@
For this mixture latent variable model (LVM), we treat $\mathcal{R}$ as latent variables; $p(\mathcal{R})$ is the prior distribution for the latent logic tress; and the joint likelihood of the event sequences $p(X,Y|\mathcal{R})$ can be derived from temporal point process framework, as shown in Eq. (5).
We will employ LLMs as the prior $p(\mathcal{R})$ for logic trees. Additionally, if we aim to leverage the powerful reasoning and generation capabilities of LLMs to predict $Y$ — for instance, in the context of symptom-treatment pairs or question-answer pairs for $(X,Y)$ — it becomes intriguing to explore the recommendations of $Y$ given $X$ provided by LLMs. Consequently, we further decompose the mixture LVM (as shown in Eq. (6)) as:
@@
$=p_{LM}(Y|X,\mathcal{R})\sum_{\mathcal{R}}p_{w}(X|\mathcal{R})p_{LM}(\mathcal{R};\phi).$ (8)
We aim to jointly optimize the event likelihood parameter $w$ and the tunable parameters $\phi$ in the prior language model. The challenge in learning arises from the latent variables $\mathcal{R}$. Fortunately, the EM algorithm provides an effective tool for learning mixture models with latent variables. However, in the E-step, we need to analytically evaluate the current posterior $p(\mathcal{R}|X,Y)\propto p_{LM}(Y|X,\mathcal{R})p_{w}(X|\mathcal{R})p_{LM}(\mathcal{R})$, which is intractable due to that the partition function requires the summation over the discrete space of $\mathcal{R}$. To tackle this intractability, variational-EM algorithm (Dempster et al., 1977; Beal, 2003; Koller & Friedman, 2009) can be used to approximate the posterior by optimization. We will address this issue by introducing an amortized EM, where in the E-step we learn GFlowNets to sample from $p(\mathcal{R}|X,Y)$ without the need to calculate the partition function.
@@ 3.2 Amortized EM framework for Logic Tree Inference
Specifically, in the E-step, we will draw samples from the posterior over the latent symbolic logic tree, denoted as $p_{LM}(\mathcal{R}|X,Y)$, which comes from an amortized sampler of $\mathcal{R}$ with an LLM as its policy. In the M-step, we maximize the log-likelihood of the joint probability of the sampled latent variables $\mathbb{E}_{\mathcal{R}\sim p(\mathcal{R}|X,Y)}[\log p_{LM}(Y|X,\mathcal{R})p_{w}(X|\mathcal{R})p_{LM}(\mathcal{R})]$ with respect to the parameters of $w$. This combination of amortized inference (learning to sample the symbolic logic tree from the language model) and supervised learning (optimizing the likelihood model with the ‘supervision’ involving $\mathcal{R}$ sampled from the amortized posterior) is presented in Fig. 2. We illustrate them in detail in the sections below.
@@
$\log q_{GFN}(\mathcal{R}_{t+1}|\mathcal{R}_{t}):=\sum_{i=1}^{m}\sum_{k=1}^{W+1}\log q_{LM}(z_{j+1}^{(i),k}|z_{0}^{(i)}\cdots z_{j}^{(i)}),$ (9)
@@
The marginal likelihood of sampling a terminal logic tree $\mathcal{R}_{t}$ is given by
$q_{GFN}(\mathcal{R}_{t}\to\mathtt{T})=$
@@
over trajectories $\tau$ starting at $\mathcal{R}_{0}$ and ends at $\mathcal{R}_{t}$. Notably, the goal of GFlowNet training is to fit the parametric policy $q_{GFN}(\cdot|\cdot;\theta)$ such that its terminating probability $q_{GFN}(\mathcal{R}_{t}\to\mathtt{T})$ is proportional to a predefined reward $r$. In our case, GFlowNet’s reward $r$ is defined as the posterior of logic trees $p(\mathcal{R}|X,Y)$, i.e., $r(\mathcal{R}|X,Y)\propto p_{LM}(Y|X,\mathcal{R})p_{w}(X|\mathcal{R})p_{LM}(\mathcal{R})$. By construction, GFlowNet’s marginal terminating distribution is proportional to its reward function $r(\mathcal{R}|X,Y)$, thus we will have the final samples $\mathcal{R}$ given by the GFlowNet’s policy $q_{GFN}(\cdot|\cdot,\theta)$ following the distribution of unnormalized true posterior of $p(\mathcal{R}|X,Y)$. Here, the given reward function $r$ can be decomposed as a product of likelihood terms that accumulated over steps of the sampling sequence. In this case, a forward-looking SubTB loss (Madan et al., 2023) for GFlowNet can help local credit assignment (Hu et al., 2023a, b). The SubTB learning objective for trajectory $\tau=(\mathcal{R}_{0},\mathcal{R}_{1},\cdots,\mathcal{R}_{t})$ is:
@@
where $q_{\theta}$ is the conditional GFlowNet policy initialized by a language model $p_{LM}$ conditioned on prefix $X$ and $Y$. The detailed derivation of the SubTB loss is given in Appendix F. In practice, this loss can be minimized by gradient descent on $\theta$ sampled either on-policy or off-policy, just as in reinforcement learning. To predict the event type $Y$ for an unseen event sequence $X$, one can draw samples of $\mathcal{R}$ from $q_{\theta}$ followed by sampling from $p_{LM}(Y|X,\mathcal{R})$.
@@
  until some convergence criteria
@@ 4.1 Experimental Setup
Competitors. In our study, we categorize competitors into three distinct types. The first category includes prompt-based approaches applied to language models, such as $k$-shot Chain-of-Thought (CoT) (Wei et al., 2022) and Tree-of-Thoughts (ToT) (Yao et al., 2023), which are utilized to generate reasoning chains and make prediction of the last event. The second category involves fine-tuning methods for language models, notably supervised fine-tuning (SFT) and Proximal Policy Optimization (PPO) (Schulman et al., 2017) fine-tuning. The final category consists of advanced neural Temporal Point Process (TPP) models specifically designed for event prediction. Within this group, we focus on AttNHP (Yang et al., 2021), an attention-based TPP whose performance is either on par with or superior to the Neural Hawkes Process (NHP) (Mei & Eisner, 2017) and other attention-based models (Xue et al., 2023a). Additionally, we consider PromptTPP (Xue et al., 2023b), a prompting model based on AttNHP (abbreviated as Pt-AttNHP), tailored for processing streaming events with a retrieval memory mechanism.
@@ Acknowledgement
The authors thank the anonymous reviewers for their careful reading of our manuscript and their many insightful comments and suggestions. Shuang Li’s research was in part supported by the National Science and Technology Major Project under grant No. 2022ZD0116004, the NSFC under grant No. 62206236, Shenzhen Science and Technology Program under grant No. JCYJ20210324120011032, Shenzhen Key Lab of Cross-Modal Cognitive Computing under grant No. ZDSYS20230626091302006, and Guangdong Key Lab of Mathematical Foundations for Artificial Intelligence. Zitao Song and Bo AN are supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-009).